A final exam in Statistics has a mean of 73 with a standard deviation of 7.73. Assume that a random sample of 24 students is selected and the mean test score of the sample is computed. What percentage of sample means are less than 70?
Answer by stanbon(75887)

A final exam in Statistics has a mean of 73 with a standard deviation of 7.73. Assume that a random sample of 24 students is selected and the mean test score of the sample is computed. What percentage of sample means are less than 70?
----
std of the sample means for samples of size 2 is 7.73/sqrt(24)
-----------
z(70) = (70-73)/[7.73/sqrt(24)] = 1.9013
----
P(x < 70) = P(z < 1.9013) = normalcdf(-100,1.9013) = 0.9714
================================
Cheers,
Stan H.
===================